OpenStudy (anonymous):

-10x+y=25 find x and y intercept

OpenStudy (anonymous):

x intercept is the point on the x axis this means that y would be equal to 0 so substitute 0 for y and then solve for x y intercept is the point on the y axis this means that x would be equal to 0 substitute 0 for x and solve for y

OpenStudy (anonymous):

let x = 0 to find y let y = 0 to find x

OpenStudy (anonymous):

-10x+0=25 -10*0+y=25

OpenStudy (anonymous):

keep going

OpenStudy (anonymous):

-10x=25 y=25

OpenStudy (anonymous):

so the y intercept would be 25 what would the x intercept be?

OpenStudy (anonymous):

I'm lost

OpenStudy (anonymous):

-10x+y=25 y intercept means x=0 so y=25 that means the y-intercept is (0,25) x intercept means y=0 -10x=25 solve for x

OpenStudy (anonymous):

still lost?

OpenStudy (anonymous):

-10x/-10=25/-10 x=25/10

OpenStudy (anonymous):

you forgot the negative sign

OpenStudy (anonymous):

x=25/-10

OpenStudy (anonymous):

and thus the x-intercept is

OpenStudy (anonymous):

5/-2=x

OpenStudy (anonymous):



OpenStudy (anonymous):

Thank you so much everyone

OpenStudy (anonymous):

yw :)